### v1

@@ IV-D Time-Invariant Limit Dynamics
The next result uses Assumptions (A6)-(A9) to show the existence and attractivity of stationary solutions, which will then be used to prove the LTI version of Theorem IV.8.
Theorem IV.11.
@@
Proof.
The proof is the same as proof of [1, Thm IV.4]. ∎
Combining the linearity of Theorem IV.8 and the stationarity of Theorem IV.11 ensures the convergence of the average population dynamics to an LTI system, as shown next.
Theorem IV.12.
@@
and $\mathbf{A}_{\infty}(t),\mathbf{B}_{\infty}(t)$ are as in (IV.8).
Proof.
By Theorem IV.11 each subsystem $\mathbf{x}_{i}(t)$ converges in distribution to the corresponding stationary solution $\mathbf{x}_{i}^{*}\sim p_{i}^{*}$ as $t\to\infty$, which are $\rho$-mixing (in the limit) by (A10). Define
@@
and let $\bar{\mathbf{x}}^{*}_{\infty}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\bar{x}^{*}})$ where $\boldsymbol{\Sigma}_{\bar{x}^{*}}={\rm Cov}(\bar{\mathbf{x}}^{*})$. Then by Theorem IV.6, $\bar{\mathbf{x}}^{*}$ converges in distribution to $\bar{\mathbf{x}}^{*}_{\infty}$ as $N\to\infty$ and, therefore,
@@
It then follows from (42) that all the time-dependent covariances in (IV.8) converge to their respective limits and, hence, so do the matrices $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$. This completes the proof. ∎
Similar to the guaranteed linearity of Theorem IV.8, the result of Theorem IV.12 is asymptotic and LTI dynamics are only approached as $N$ and $t$ grow to infinity. In many real-world systems, however, dynamics settle to stationary solutions after only a few time steps. Further, and more importantly for our discussion, nonlinearities often vanish rather quickly with averaging. We next focus on formalizing the latter dimension, namely, the rate of convergence to linearity under spatial averaging.
IV-E Rate of Convergence to Linearity
In this section we seek to estimate the rate at which spatially-averaged dynamics converge to linearity. We have empirically shown that this rate is often rather fast [1], but rigorous characterizations are missing. The main result of this section is the following, where we prove that convergence of conditional expectations (similar to (IV.8)) occurs at the $O(1/\sqrt{N})$ rate. We present this result for the simplest case of i.i.d. sequences of static random vectors and discuss generalizations afterwards.
Theorem IV.13.
@@ IV-E Rate of Convergence to Linearity
where in the latter inequality we used $\big|\frac{a}{b}-\frac{c}{d}\big|\leq\frac{|d||a-c|+|c||b-d|}{|b||d|}$. Now, using the standard relationship between $\mathbb{E}[X]=-\int_{0}^{\infty}F_{X}(x)dx+\int_{0}^{\infty}(1-F_{X}(x))dx$ between the cumulative distribution function (CDF) and expected value of any random variable, we can translate the bound on the difference of conditional CDFs in (IV-E) into a bound on the difference of conditional expectations, as follows.
@@
Note that, because the bound in (IV-E) is independent of $\xi$, substituting it directly into (IV-E) will give a useless bound of $\infty$. However, because $x_{i}\in[-M,M]$ by assumption, $F_{\bar{x}_{i}|\bar{\mathbf{y}}}\big(\xi|\mathcal{B}_{\delta}(\boldsymbol{\upsilon})\big)=0$ for all $\xi\notin[-M,M]$ and we can use this to achieve a finite bound, as follows. For reasons that we will see later, let $M^{\prime}\geq M$ be a (yet-to-be-determined) more conservative bound on $x_{i}$. Then,
@@
To evaluate the integral in (IV-E), note that
@@
Notice that, from (IV-E), always $\xi\leq-M^{\prime}$, where $M^{\prime}\geq M$ is yet to be determined. Choose $M^{\prime}=\max\{M,1-\xi_{0}\}$. Then, for all $\xi\leq-M^{\prime}$ we have
@@
Substituting this into (IV-E) and integrating from -$\infty$ to $M^{\prime}$ gives
@@
Because this is true for all $i=1,\dots,n$, and $\|\cdot\|\leq\sqrt{n}\|\cdot\|_{\infty}$, we get (44), completing the proof. ∎
@@
A number of remarks about Theorem IV.13 are in order. First, this is the only result in this paper where we assume independence among the subsystems, in order to simplify the analysis. If the subsystems form a $\rho$-mixing sequence, similar results can be obtained using extensions of the Berry-Esseen theorem (see, e.g., [45]) but the analysis becomes more involved and the rates of convergence become variable depending on the rate at which correlations decay (e.g., $n^{-\frac{1}{6}}\log n$ if correlations decay as $n^{-\frac{3}{2}}$). Second, the $L_{\infty}$ nature of the Berry-Esseen theorem (i.e., that the bound in (45) does not depend on the set $U$) has forced us to move from the singleton-conditioning in (IV.8) to neighborhood-conditioning in (44). To the best of our knowledge this is not avoidable for continuous-valued conditioning variables (see [46] for conditioning on discrete variables) but the radius of $\mathcal{B}_{\delta}(\boldsymbol{\upsilon})$ can be made sufficiently small so that conditioning on $\bar{\mathbf{y}}\in\mathcal{B}_{\delta}(\boldsymbol{\upsilon})$ is a good approximation of conditioning on $\bar{\mathbf{y}}=\boldsymbol{\upsilon}$. Finally, we presented Theorem IV.13 for static random variables, but its extension to dynamical systems is straightforward, with $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ in (43) being replaced by $\mathbf{x}_{i}(t+1)$ and $\mathbf{y}_{i}(t)=\begin{bmatrix}\mathbf{x}_{i}(t)^{T}&\mathbf{w}_{i}(t)^{T}\end{bmatrix}^{T}$, respectively. Under the same assumptions, namely, i.i.d. subsystems with uniformly bounded states and noises, the same conclusion as in (44) would then hold for conditional expectation of average states at any time $t+1$ given neighborhood (not singleton) conditions on average states and noises at the previous time $t$.
@@
V Linearizing Effect of Spatial Averaging on Spatially-Embedded Dynamical Systems
V-A Problem Formulation
@@
While the results of Section IV are general in many respects (form of nonlinearity, noise characteristics, etc.) they are still limited in requiring a decay of correlations with the difference in two subsystems’ linear indices (i.e., $\varrho(|i-j|)\to 0$ as $|i-j|\to\infty$). In brain networks that initially motivated the present work [19], e.g., it is well-known that correlations between neurons decay as their distance grows [47, 48, 49], but this distance is physical (Euclidean) rather than between linear indices. Similarly, common forms of spatial averaging that occur in the brain, such as those underlying functional magnetic resonance imaging (fMRI) or electroencephalography (EEG), occur over neurons that are close to each other in a physical sense rather than index-wise. Therefore, in this section we generalize our results of Section IV to systems where both the decay of pairwise correlations and the domains of spatial averaging occur in Euclidean space.
@@
The rest of this section is concerned with addressing Problem 2. We will follow the same general steps as in Section IV, focusing on what needs to be done differently in order to address the challenges that arise from the new formulation.
V-B $\rho^{*}$-Mixing Sequences
@@
$\rho^{*}$-mixing sequences share many of the same properties with $\rho$-mixing ones. For example, the notion of a residual factor (Definition IV.2) extends naturally to $\rho^{*}$-mixing sequences. The formal definition, for clarity, is as follows.
Definition V.2.
@@
In what follows we also need the following result, which parallels Lemma IV.5 and can be proved using the same approach.
Lemma V.3.
@@
Next we move to the CLT, which lies at the core of our methodology. Here the main challenge lies in proving a result similar to Proposition III.3 for $\rho^{*}$-mixing sequences, as tackled next. Once that is shown, generalization to multivariate sequences will be straightforward.
Theorem V.4.
@@
where in the last equality we used the fact that $\lim_{\mathcal{R}\to\infty}h\big(N_{\gamma^{1/d}\mathcal{R}}\big)/h(N_{\mathcal{R}})=1$ by the uniform convergence of slowly-varying functions [50, Thm. 1.2.1].
@@
Furthermore, from [51, Prop 3.11], for any two $\sigma$-algebras $\mathcal{A}$ and $\mathcal{B}$,
@@
which is the same as (V-B).
Put together, it follows from (60), (V-B), (62), (63), and (V-B) that $W_{\mathcal{R}}(\gamma)$ satisfies all the assumptions of [52, Thm 19.2], and hence
@@
which completes the proof. ∎
A minor difference between Theorem V.4 and Proposition III.3 is the difference between assumptions (58b) and (4b). The latter are both technical assumptions needed to prove uniform integrability of families of partial sum processes such as that in (62), which in turn prevents pathological cases where some probability mass gradually drifts to infinity as $N$ or $\mathcal{R}\to\infty$. Both assumptions are sufficient but not necessary for uniform integrability, neither is uniformly stronger or weaker than the other, and either can be replaced by the direct (but perhaps more opaque) assumption of uniform integrability.
Similar to Section IV, we next generalize Theorem V.4 to multivariate sequences. The proof technique is the same as that of Theorem IV.6 and hence omitted.
Theorem V.5.
@@
V-C Asymptotic Linearity Under Spatial Averaging
Theorem V.5 provides the basis that we need for solving Problem 2. However, note that we have so far assumed, for simplicity, that the Euclidean locations $\{\mathbf{r}_{i}\}_{i=1}^{\infty}$ are deterministic and fixed–e.g., at the vertices of a lattice–whereas in reality spatially-embedded subsystems are often distributed randomly over space. Therefore, before getting to the main result of this section, we prove the following lemma showing that our only requirement on the subsystems’ locations, i.e., assumption (58c), is still satisfied with probability 1 in a proper stochastic setting.
Lemma V.6.
@@ V-C Asymptotic Linearity Under Spatial Averaging
and let $|A_{\ell}|$ be the volume of $A_{\ell}$. Because $A_{\ell}$’s are disjoint, by definition [53, Def. 3.1],
@@
The statement of the theorem then follows by replacing $r=C_{v}\mathcal{R}(r)^{d}$ in (71) and changing the limiting variable from $r$ to $\mathcal{R}$. ∎
Lemma V.6 allows us to restate the preceding CLTs for $\rho^{*}$-mixing sequences with fixed Euclidean locations for the case where the locations are Poisson distributed. These restatements are straightforward and hence omitted. Instead, we proceed to use the combination of Lemma V.6 and the preceding CLTs to solve Problem 2. The following is the parallel to Assumption IV.7 for spatially-embedded systems.
Assumption V.7.
@@
We are now ready to state the main result of this section, as follows. The proof is similar to the proof of Theorem IV.8 and hence omitted.
Theorem V.8.
@@
where $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$ are defined in (IV.8). Furthermore, if the subsystem locations $\{\mathbf{r}_{i}\}_{i=1}^{\infty}$ are distributed according to a homogeneous Poisson point process on $\mathbb{R}^{d}$, (72) holds with probability 1. $\square$
As we noted in the Introduction,the present work was motivated by our empirical observations of macroscopic linearity in brain networks. Theorem V.8 now rigorously explains these observations, and highlights the key roles of spatial averaging that is inherent in macroscopic recordings such as electroencephalography (EEG) [54] and functional magnetic resonance imaging (fMRI) [55], as well as the weak and decaying correlations that have long been observed between biological neurons [47, 48, 49]. Whether and to what extent large-scale complex networks in other domains meet the assumptions of this theorem, and are hence bound to follow its consequent macroscopic linearity, remains an interesting topic of inquiry for future research.
VI Conclusions

### v2

@@ 4.4 Time-Invariant Limit Dynamics
The next result uses Assumptions (A6)-(A9) to show the existence and attractivity of stationary solutions, which will then be used to prove the LTI version of Theorem 4.8.
Theorem 4.11.
@@
{proof}
The proof is the same as proof of [1, Thm IV.4].
Combining the linearity of Theorem 4.8 and the stationarity of Theorem 4.11 ensures the convergence of the average population dynamics to an LTI system, as shown next.
Theorem 4.12.
@@
and $\mathbf{A}_{\infty}(t),\mathbf{B}_{\infty}(t)$ are as in (4.8).
{proof}
By Theorem 4.11 each subsystem $\mathbf{x}_{i}(t)$ converges in distribution to the corresponding stationary solution $\mathbf{x}_{i}^{*}\sim p_{i}^{*}$ as $t\to\infty$, which are $\rho$-mixing (in the limit) by (A10). Define
@@
and let $\bar{\mathbf{x}}^{*}_{\infty}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\bar{x}^{*}})$ where $\boldsymbol{\Sigma}_{\bar{x}^{*}}={\rm Cov}(\bar{\mathbf{x}}^{*})$. Then by Theorem 4.6, $\bar{\mathbf{x}}^{*}$ converges in distribution to $\bar{\mathbf{x}}^{*}_{\infty}$ as $N\to\infty$ and, therefore,
@@
It then follows from (42) that all the time-dependent covariances in (4.8) converge to their respective limits and, hence, so do the matrices $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$. This completes the proof.
Similar to the guaranteed linearity of Theorem 4.8, the result of Theorem 4.12 is asymptotic and LTI dynamics are only approached as $N$ and $t$ grow to infinity. In many real-world systems, however, dynamics settle to stationary solutions after only a few time steps. Further, and more importantly for our discussion, nonlinearities often vanish rather quickly with averaging. We next focus on formalizing the latter dimension, namely, the rate of convergence to linearity under spatial averaging.
4.5 Rate of Convergence to Linearity
In this section we seek to estimate the rate at which spatially-averaged dynamics converge to linearity. We have empirically shown that this rate is often rather fast [1], but rigorous characterizations are missing. The main result of this section is the following, where we prove that convergence of conditional expectations (similar to (4.8)) occurs at the $O(1/\sqrt{N})$ rate. We present this result for the simplest case of i.i.d. sequences of static random vectors and discuss generalizations afterwards.
Theorem 4.13.
@@ 4.5 Rate of Convergence to Linearity
where in the latter inequality we used $\big|\frac{a}{b}-\frac{c}{d}\big|\leq\frac{|d||a-c|+|c||b-d|}{|b||d|}$. Now, using the standard relationship between $\mathbb{E}[X]=-\int_{0}^{\infty}F_{X}(x)dx+\int_{0}^{\infty}(1-F_{X}(x))dx$ between the cumulative distribution function (CDF) and expected value of any random variable, we can translate the bound on the difference of conditional CDFs in (4.5) into a bound on the difference of conditional expectations, as follows.
@@
Note that, because the bound in (4.5) is independent of $\xi$, substituting it directly into (4.5) will give a useless bound of $\infty$. However, because $x_{i}\in[-M,M]$ by assumption, $F_{\bar{x}_{i}|\bar{\mathbf{y}}}\big(\xi|\mathcal{B}_{\delta}(\boldsymbol{\upsilon})\big)=0$ for all $\xi\notin[-M,M]$ and we can use this to achieve a finite bound, as follows. For reasons that we will see later, let $M^{\prime}\geq M$ be a (yet-to-be-determined) more conservative bound on $x_{i}$. Then,
@@
To evaluate the integral in (4.5), note that
@@
Notice that, from (4.5), always $\xi\leq-M^{\prime}$, where $M^{\prime}\geq M$ is yet to be determined. Choose $M^{\prime}=\max\{M,1-\xi_{0}\}$. Then, for all $\xi\leq-M^{\prime}$ we have
@@
Substituting this into (4.5) and integrating from -$\infty$ to $M^{\prime}$ gives
@@
Because this is true for all $i=1,\dots,n$, and $\|\cdot\|\leq\sqrt{n}\|\cdot\|_{\infty}$, we get (44), completing the proof.
@@
A number of remarks about Theorem 4.13 are in order. First, this is the only result in this paper where we assume independence among the subsystems, in order to simplify the analysis. If the subsystems form a $\rho$-mixing sequence, similar results can be obtained using extensions of the Berry-Esseen theorem (see, e.g., [50]) but the analysis becomes more involved and the rates of convergence become variable depending on the rate at which correlations decay (e.g., $n^{-\frac{1}{6}}\log n$ if correlations decay as $n^{-\frac{3}{2}}$). Second, the $L_{\infty}$ nature of the Berry-Esseen theorem (i.e., that the bound in (45) does not depend on the set $U$) has forced us to move from the singleton-conditioning in (4.8) to neighborhood-conditioning in (44). To the best of our knowledge this is not avoidable for continuous-valued conditioning variables (see [16] for conditioning on discrete variables) but the radius of $\mathcal{B}_{\delta}(\boldsymbol{\upsilon})$ can be made sufficiently small so that conditioning on $\bar{\mathbf{y}}\in\mathcal{B}_{\delta}(\boldsymbol{\upsilon})$ is a good approximation of conditioning on $\bar{\mathbf{y}}=\boldsymbol{\upsilon}$. Finally, we presented Theorem 4.13 for static random variables, but its extension to dynamical systems is straightforward, with $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ in (43) being replaced by $\mathbf{x}_{i}(t+1)$ and $\mathbf{y}_{i}(t)=\begin{bmatrix}\mathbf{x}_{i}(t)^{T}&\mathbf{w}_{i}(t)^{T}\end{bmatrix}^{T}$, respectively. Under the same assumptions, namely, i.i.d. subsystems with uniformly bounded states and noises, the same conclusion as in (44) would then hold for conditional expectation of average states at any time $t+1$ given neighborhood (not singleton) conditions on average states and noises at the previous time $t$.
@@
5 Linearizing Effect of Spatial Averaging on Spatially-Embedded Dynamical Systems
5.1 Problem Formulation
@@
While the results of Section 4 are general in many respects (form of nonlinearity, noise characteristics, etc.) they are still limited in requiring a decay of correlations with the difference in two subsystems’ linear indices (i.e., $\varrho(|i-j|)\to 0$ as $|i-j|\to\infty$). In brain networks that initially motivated the present work [42], e.g., it is well-known that correlations between neurons decay as their distance grows [48, 49, 44], but this distance is physical (Euclidean) rather than between linear indices. Similarly, common forms of spatial averaging that occur in the brain, such as those underlying functional magnetic resonance imaging (fMRI) or electroencephalography (EEG), occur over neurons that are close to each other in a physical sense rather than index-wise. Therefore, in this section we generalize our results of Section 4 to systems where both the decay of pairwise correlations and the domains of spatial averaging occur in Euclidean space.
@@
The rest of this section is concerned with addressing Problem 2. We will follow the same general steps as in Section 4, focusing on what needs to be done differently in order to address the challenges that arise from the new formulation.
5.2 $\rho^{*}$-Mixing Sequences
@@
$\rho^{*}$-mixing sequences share many of the same properties with $\rho$-mixing ones. For example, the notion of a residual factor (Definition 4.2) extends naturally to $\rho^{*}$-mixing sequences. The formal definition, for clarity, is as follows.
Definition 5.2.
@@
In what follows we also need the following result, which parallels Lemma 4.5 and can be proved using the same approach.
Lemma 5.3.
@@
Next we move to the CLT, which lies at the core of our methodology. Here the main challenge lies in proving a result similar to Proposition 3.3 for $\rho^{*}$-mixing sequences, as tackled next. Once that is shown, generalization to multivariate sequences will be straightforward.
Theorem 5.4.
@@
where in the last equality we used the fact that $\lim_{\mathcal{R}\to\infty}h\big(N_{\gamma^{1/d}\mathcal{R}}\big)/h(N_{\mathcal{R}})=1$ by the uniform convergence of slowly-varying functions [5, Thm. 1.2.1].
@@
Furthermore, from [7, Prop 3.11], for any two $\sigma$-algebras $\mathcal{A}$ and $\mathcal{B}$,
@@
which is the same as (5.2).
Put together, it follows from (60), (5.2), (62), (63), and (5.2) that $W_{\mathcal{R}}(\gamma)$ satisfies all the assumptions of [4, Thm 19.2], and hence
@@
which completes the proof.
A minor difference between Theorem 5.4 and Proposition 3.3 is the difference between assumptions (58b) and (4b). The latter are both technical assumptions needed to prove uniform integrability of families of partial sum processes such as that in (62), which in turn prevents pathological cases where some probability mass gradually drifts to infinity as $N$ or $\mathcal{R}\to\infty$. Both assumptions are sufficient but not necessary for uniform integrability, neither is uniformly stronger or weaker than the other, and either can be replaced by the direct (but perhaps more opaque) assumption of uniform integrability.
Similar to Section 4, we next generalize Theorem 5.4 to multivariate sequences. The proof technique is the same as that of Theorem 4.6 and hence omitted.
Theorem $5.$
@@
5.3 Asymptotic Linearity Under Spatial Averaging
Theorem 5.5 provides the basis that we need for solving Problem 2. However, note that we have so far assumed, for simplicity, that the Euclidean locations $\{\mathbf{r}_{i}\}_{i=1}^{\infty}$ are deterministic and fixed–e.g., at the vertices of a lattice–whereas in reality spatially-embedded subsystems are often distributed randomly over space. Therefore, before getting to the main result of this section, we prove the following lemma showing that our only requirement on the subsystems’ locations, i.e., assumption (58c), is still satisfied with probability 1 in a proper stochastic setting.
Lemma 5.6.
@@ 5.3 Asymptotic Linearity Under Spatial Averaging
and let $|A_{\ell}|$ be the volume of $A_{\ell}$. Because $A_{\ell}$’s are disjoint, by definition [33, Def. 3.1],
@@
The statement of the theorem then follows by replacing $r=C_{v}\mathcal{R}(r)^{d}$ in (71) and changing the limiting variable from $r$ to $\mathcal{R}$.
Lemma 5.6 allows us to restate the preceding CLTs for $\rho^{*}$-mixing sequences with fixed Euclidean locations for the case where the locations are Poisson distributed. These restatements are straightforward and hence omitted. Instead, we proceed to use the combination of Lemma 5.6 and the preceding CLTs to solve Problem 2. The following is the parallel to Assumption 4.7 for spatially-embedded systems.
Assumption 5.7.
@@
We are now ready to state the main result of this section, as follows. The proof is similar to the proof of Theorem 4.8 and hence omitted.
Theorem 5.8.
@@
where $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$ are defined in (4.8). Furthermore, if the subsystem locations $\{\mathbf{r}_{i}\}_{i=1}^{\infty}$ are distributed according to a homogeneous Poisson point process on $\mathbb{R}^{d}$, (72) holds with probability 1. $\square$
As we noted in the Introduction,the present work was motivated by our empirical observations of macroscopic linearity in brain networks. Theorem 5.8 now rigorously explains these observations, and highlights the key roles of spatial averaging that is inherent in macroscopic recordings such as electroencephalography (EEG) [14] and functional magnetic resonance imaging (fMRI) [13], as well as the weak and decaying correlations that have long been observed between biological neurons [48, 49, 44]. Whether and to what extent large-scale complex networks in other domains meet the assumptions of this theorem, and are hence bound to follow its consequent macroscopic linearity, remains an interesting topic of inquiry for future research.
6 Conclusions